8 children divide 32 apples. Kitty took 1 apple, Mary 2, Sherrie 3, and Caroline 4. Bob Smith took as many as his sister. Ken Brown took twice as many as his sister. Jim Jones 3 times as many as his sister and Tom Robinson 4 times as many as his sister. What is the last name of each of the 4 girls?
Please help!! I think I have tried every combo but the closest I have come is a total of 29. Thanks.

Answer by ankor@dixie-net.com(12701)

8 children divide 32 apples. Kitty took 1 apple, Mary 2, Sherrie 3, and Caroline 4. Bob Smith took as many as his sister. Ken Brown took twice as many as his sister. Jim Jones 3 times as many as his sister and Tom Robinson 4 times as many as his sister. What is the last name of each of the 4 girls?
:
I came up with Kitty Jones, Mary Robinson, Sherrie Smith, and Caroline Brown
K: 1*3 + 1 = 4
M: 2*4 + 2 = 10
S: 3*1 + 3 = 6
C: 4*2 + 4 = 12
----------------
for a total 32
:
However, I can't come up with a neat algebra way to doing it.
I reasoned that the average total of each of the 4 pairs is 8
Therefore 2 pair totals will be above 8 and 2 pair totals below 8 but that
did not really help much